OpenStudy (anonymous):

Suppose that in 2001, the average cost of a gallon of milk was $2.43. In 2007 the average price was $4.21. Assuming the relationship is linear, find the equation of the line that models the price of milk over time since 1999. Round to the nearest thousandth.

OpenStudy (anonymous):

4.21 = 2.43*(1 + i)^6 i = 0.096*100 = 9.6% is the increase on the price of milk in the considered period

OpenStudy (anonymous):

HINT: use two point formula